OpenStudy (anonymous):

please help! Two angles are supplementary. Three times the measure of the smaller is 4 degrees less than the measure of the larger. true or false?

OpenStudy (anonymous):

L+S=180 and 3S = L-4 now solve and see for yourself..

OpenStudy (anonymous):

i got s=180-L 3(180-L) 540-3L=L-4 -4L=-544 L=136 then i pluged in to the first equation and i got s=44 but what does that mean? :/

OpenStudy (anonymous):

that means the given case is possible and true.. right?

OpenStudy (anonymous):

uhmm.... sure! but i'm still a little confused can you explain further? how did you know it was true?

OpenStudy (anonymous):

see if it wasn't true.. you won't get a value that satisfies the given condition.. Like if the given statement were false, you would get the values of angles like in -ve quantity or larger than 180... (both of which would be false)...

OpenStudy (anonymous):

ohhh i see! thanks!